Reports 1-1 of 1 Clear search Modify search
MIF (ITF Control)
dan.chen - 17:09 Tuesday 18 June 2024 (29964) Print this report
Comment to MICH lock trial (29949)

With Yuzurihara, Tanaka, Yokozawa, YamaT, Ushiba

Summary

The BS alignment was adjusted following the ITMY alignment, with final P and Y values of (40.7, -17.4) and a measured contrast of 92.2%.
A 6dB gain difference was investigated, initially suspecting laser power loss and clipping, but measurements showed no clipping and significant reductions in laser power across multiple sensors.
The actual issue was incorrect phase adjustment on POP_PDA1_RF17. Correcting the phase adjustment resulted in only a minor gain difference of 0.8 dB.

Details

  1. Because ITMY alignment was adjusted in this morning as reported on [klog29952](https://klog.icrr.u-tokyo.ac.jp/osl/?r=29952), we adjusted the alignment of BS.

    1. `K1:VIS-BS_TM_SET_P_OFFSET` and `K1:VIS-BS_TM_SET_P_OFFSET` were adjusted by moniting POP PDA1 `K1-LSC-POP_PDA1_DC_OUT_DQ`.

    2. The final P and Y values: (40.7, -17.4)

    3. Measured the contrast: 92.2% (fig1)

  2. Investigation of 6dB difference in gain from the previous measurement (Measured at 01/11/2022 10:06:50)

    1. Lock conditions

      1. VERTEX: MICH_LOCKED state

      2. VIS_ITMY: ALIGNED (This is to prevent oscillation of ITMY)

      3. Filters in K1LSC-MICH2: ELP300=ON, OBS=OFF

    2. Suspect loss of laser power

      1. Check the laser power (fig2)

      2. Clipping survey

        1. Changed state: VERTEX=DOWN, ITMX=LOCK_ACQUISITION, ITMY=MISALINED

        2. We changed P,Y in IMMT1 (`K1:VIS-IMMT1_TM_TEST_P_OFFSET` and `K1:VIS-IMMT1_TM_TEST_Y_OFFSET`) and looked at POP PDA1 and REFL PDA1 to see if the signal was reduced.

        3. Result

          • TEST_P_OFFSET = 0 -> REFL_PDA1 = 0.08

          • TEST_P_OFFSET = +18000 -> REFL_PDA1 = 0.076

          • TEST_P_OFFSET = -18000 -> REFL_PDA1 = 0.04

          • TEST_Y_OFFSET = 0 -> REFL_PDA1 = 0.08

          • TEST_Y_OFFSET = 3000 -> REFL_PDA1 = 0.074

          • TEST_Y_OFFSET = -3000 -> REFL_PDA1 = 0.058

          • (POP_PDA1 did not change at any case.)

        4. We also performed the same task as IMMT1.

        5. From thses results, we belive the beams are not clipping.

    3. We checked phasing of POP_PDA1_RF17 as reported on [klog29961](https://klog.icrr.u-tokyo.ac.jp/osl/?r=29961)

      1. As reported on the klog, we used MICH to adjusted the phase, which makes produced different results.

      2. After the phase adjustment, we locked MICH with VERTEX guardian like what we did above.

  3. OLTF measurement again

    1. Meausred result is the following. (fig3)

    2. Note

      • This measured result(red on the graph)

      • As for GAIN, the present results (red) were only 0.8 dB larger, but they were almost identical. Accuracy of phasing made this diff?

      • Phase are different from the old measurement because the frequency of ELP is defferent (old=100Hz, current=300Hz)

      • In order to have a good comparizon, we applied 1.14e-9 to the current measurement, which containns the unit different of MICH1_IN1 and the gain different of MICHcal in POP_PDA1_RF17 filter banks.

Images attached to this comment
Search Help
×

Warning

×